# Content-based image retrieval for multi-class volumetric radiology images: a benchmark study

### v1

@@ Abstract
While content-based image retrieval (CBIR) has been extensively studied in natural image retrieval, its application to medical images presents ongoing challenges, primarily due to the 3D nature of medical images. Recent studies have shown the potential use of pre-trained vision embeddings for CBIR in the context of radiology image retrieval. However, a benchmark for the retrieval of 3D volumetric medical images is still lacking, hindering the ability to objectively evaluate and compare the efficiency of proposed CBIR approaches in medical imaging. In this study, we extend previous work and establish a benchmark for region-based and multi-organ retrieval using the TotalSegmentator dataset (TS) with detailed multi-organ annotations. We benchmark embeddings derived from pre-trained supervised models on medical images against embeddings derived from pre-trained unsupervised models on non-medical images for 29 coarse and 104 detailed anatomical structures in volume and region levels. We adopt a late interaction re-ranking method inspired by text matching for image retrieval, comparing it against the original method proposed for volume and region retrieval achieving retrieval recall of 1.0 for diverse anatomical regions with a wide size range. The findings and methodologies presented in this paper provide essential insights and benchmarks for the development and evaluation of CBIR approaches in the context of medical imaging.
@@ 1 Introduction
In the realm of computer vision, content-based image retrieval (CBIR) has been the subject of extensive research for several decades (Dubey, 2021). CBIR systems commonly preserve low-dimensional image representations in a database and subsequently retrieve similar images based on the distance/similarity of the image representations (Khun Jush et al., 2023). Early approaches to CBIR involved manually crafting distinctive features, which led to a semantic gap, resulting in the loss of crucial image details due to the limitations of low-dimensional feature design (Dubey, 2021; Wang et al., 2022). However, recent studies in deep learning have redirected attention towards the creation of machine-generated discriminative feature spaces, effectively addressing and bridging this semantic gap (Qayyum et al., 2017). This shift has significantly enhanced the potential for more accurate and efficient CBIR methods (Dubey, 2021).
@@
While natural image retrieval has been extensively researched, the application of retrieval frameworks to medical images, particularly radiology images, presents ongoing challenges. CBIR offers numerous advantages for medical images. Radiologists can utilize CBIR to search for similar cases, enabling them to review the history, reports, patient diagnoses, and prognoses, thereby enhancing their decision-making process. In real-world use-cases, we often encounter huge unannotated datasets available from different studies where the DICOM headers are removed or incorrect. Finding relevant images in such databases is extremely time-consuming. Moreover, the development of new tools and research in the medical field requires trustable dataset sources and therefore a reliable method for retrieving images, making CBIR an essential component in advancing computer-aided medical image analysis and diagnosis. The challenge with applying CBIR to medical images lies in the fact that algorithms developed for natural images are typically designed for 2D images, while medical images are often 3D volumes which adds a layer of complexity to the retrieval process.
@@
Recent studies have proposed and demonstrated the potential use of pre-trained vision embeddings for CBIR in the context of radiology image retrieval (Khun Jush et al., 2023; Abacha et al., 2023; Denner et al., 2024; Truong et al., 2023). However, these studies have primarily focused on 2D images (Denner et al., 2024) or specific pathologies or tasks (Abacha et al., 2023; Khun Jush et al., 2023; Truong et al., 2023), overlooking the presence of multiple organs in the volumetric images, which is a critical aspect of real-world scenarios. Leveraging multilabel datasets can thoroughly evaluate the efficacy of the proposed methods, enabling a more comprehensive assessment of CBIR approaches for radiology images. Despite previous efforts, there is still no established benchmark available for comparing methods for the retrieval of 3D volumetric medical images. This absence of a benchmark impedes the ability to objectively evaluate and compare the efficiency of the proposed CBIR approaches in the context of medical imaging.
@@
Our previous work (Khun Jush et al., 2023) demonstrated the potential of utilizing pre-trained embeddings, originally trained on natural images, for various medical image retrieval tasks using the Medical Segmentation Decathlon Challenge (MSD) dataset (Antonelli et al., 2022). The approach is outlined in Figure 1. Building upon this, the current study extends the methodology proposed in Khun Jush et al. (2023) to establish a benchmark for anatomical region-based and multi-organ retrieval. While the focus of Khun Jush et al. (2023) was on evaluating sampling methods within the context of the single-organ MSD dataset (Antonelli et al., 2022), it was observed that the single-organ labeling, hinders the evaluations for images containing multiple organs. The main objective of this study is to set a benchmark for organ retrieval at the subvolume or region-based level, which is particularly valuable in practical scenarios, such as when users zoom in on specific regions of interest to retrieve similar images of the precise organ under examination. To achieve this, we evaluate a count-based method in regions using the TotalSegmentator dataset (TS) (Wasserthal et al., 2023). TS dataset along with its detailed multi-organ annotations is a valuable resource for medical image analysis and research. This dataset provides comprehensive annotations for 104 organs or anatomical structures, which allow us to derive fine-grained retrieval tasks and comprehensively evaluate the proposed methods.
@@
  • We benchmarked pre-trained embeddings trained supervised on medical images against pre-trained embeddings trained unsupervised on non-medical images for 29 modified coarse anatomical regions and 104 original anatomical regions from TS dataset Wang et al. (2022).
  • We adopted a late interaction re-ranking method originally used for text retrieval called ColBERT (Khattab and Zaharia, 2020) for image retrieval that takes the similarity over the whole volumes into account.
  • We benchmarked the re-ranking method against the original method proposed in Khun Jush et al. (2023) for volume and region retrieval on 29 modified coarse anatomical regions and 104 original anatomical regions from TS dataset Wang et al. (2022).
@@ 2.1 Vector Database and Indexing
In the context of image search the database is where all the representations of the images, a.k.a. embeddings, and their metadata including annotations are stored. A query allows the user or the system to request specific images in various ways, e.g., by inputting a reference image or a textual description. The goal is to search the database for similar images that match the query. Similarly, in this study, the search process entails comparing a query image with images in the database to identify the most similar image using the similarity of the embeddings. Throughout this process, we do not depend on any metadata information at any stage. Metadata-independence is an intended design choice and in stark contrast to widely used metadata-based image retrieval solutions that frequently lack the necessary specificity in real-world retrieval applications. In small sets, the similarity search is easy but with the growing size of the database, the complexity increases. Accuracy and speed are the key factors in search, thus, naive approaches typically fail in huge datasets.
@@
Indexing in the context of content-based image search involves creating a structured system that allows for efficient storage and retrieval of images based on their visual content. A flat index is the simplest form of indexing, where no modification is made to the vectors before they are used for search. In flat indexing, the query vector is compared to every other full-size vector in the database and their distances are calculated. The nearest k of the searched spaces is then returned as the k-nearest neighbors (kNN). While this method is the most accurate, it comes at the cost of significant search time Aumüller et al. (2020). To improve search time, two approaches can be employed: reducing the vector size through dimensionality reduction, e.g., by reducing the number of bits representing each vector, or reducing the search scope by clustering or organizing vectors into tree structures based on similarity or distance. This results in the identification of an approximation of the true nearest neighbors, known as approximate nearest neighbor search (ANN) (Aumüller et al., 2020).
@@
There are several ANN methods available. Khun Jush et al. (2023) compared Locality Sensitive Hashing (LSH) and Hierarchical Navigable Small World (HNSW) for indexing and search. LSH hashes data points in a way that similar data points are mapped to the same buckets with higher probabilities. This allows for a more efficient search for nearest neighbors by reducing the number of candidates to be examined. HNSW (Malkov and Yashunin, 2018) indexing organizes data into a hierarchical graph structure where each layer of the hierarchy has a lower resolution. The top layer connects data points directly, but the lower layers have fewer connections. The graph structure is designed to allow for efficient navigation during the search. Compared to LSH, HNSW enables faster search and requires less memory Taha et al. (2024). Based on findings in (Khun Jush et al., 2023) HSNW was chosen in this setting over LSH as the preferred indexing method due to speed advantages at a comparable recall.
We propose a system similar to Khun Jush et al. (2023) and Truong et al. (2023) that allows the pre-computation of the image representations of the database. There are various index solutions available to store and search vectors. In this study, we used the Facebook AI Similarity Search (FAISS) package that enables fast similarity search (Johnson et al., 2019). The indexing process involves running the feature extractors on slices of each volumetric image and storing the output embeddings per slice. The produced representations are then added to the search index which is used later on for vector-similarity-based retrieval.
@@ 2.2 Feature Extractors
Self-supervised Models: We employed three self-supervised models pre-trained on ImageNet (Deng et al., 2009). DINOv1 (Caron et al., 2021), that demonstrated learning efficient image representations from unlabeled data using self-distillation. DINOv2 (Oquab et al., 2023), is built upon DINOv1 (Caron et al., 2021), this model scales the pre-training process by combining an improved training dataset, patchwise objectives during training and introducing a new regularization technique, achieving superior performance on segmentation tasks. DreamSim (Fu et al., 2023), built upon the foundation of DINOv1 (Caron et al., 2021), fine-tunes the model using synthetic data triplets specifically designed to be cognitively impenetrable with human judgments.
Supervised Models: We included a SwinTransformer model (Liu et al., 2021) and a ResNet50 model (He et al., 2016) trained in a supervised manner using the RadImageNet dataset (Mei et al., 2022) that includes 5 million annotated CT, MRI, and ultrasound images of musculoskeletal, neurologic, oncologic, gastrointestinal, endocrine, and pulmonary pathology. Furthermore, a ResNet50 model pre-trained on rendered images of fractal geometries was included based on (Kataoka et al., 2022). These training images are formular-derived, non-natural, and do not require any human annotation.
2.3 Dataset
@@
adrenal gland left adrenal gland iliopsoas right iliopsoas rib right 8 rib
adrenal gland right adrenal gland inferior vena cava cardiovascular system rib right 9 rib
@@
gallbladder gallbladder rib left 10 rib vertebrae C5 vertebrae
gluteus maximus left gluteus muscles rib left 11 rib vertebrae C6 vertebrae
gluteus maximus right gluteus muscles rib left 12 rib vertebrae C7 vertebrae
gluteus medius left gluteus muscles rib left 2 rib vertebrae L1 vertebrae
gluteus medius right gluteus muscles rib left 3 rib vertebrae L2 vertebrae
gluteus minimus left gluteus muscles rib left 4 rib vertebrae L3 vertebrae
gluteus minimus right gluteus muscles rib left 5 rib vertebrae L4 vertebrae
heart atrium left cardiovascular system rib left 6 rib vertebrae L5 vertebrae
heart atrium right cardiovascular system rib left 7 rib vertebrae T1 vertebrae
heart myocardium cardiovascular system rib left 8 rib vertebrae T10 vertebrae
heart ventricle left cardiovascular system rib left 9 rib vertebrae T11 vertebrae
heart ventricle right cardiovascular system rib right 1 rib vertebrae T12 vertebrae
hip left hip rib right 10 rib vertebrae T2 vertebrae
hip right hip rib right 11 rib vertebrae T3 vertebrae
humerus left humerus rib right 12 rib vertebrae T4 vertebrae
humerus right humerus rib right 2 rib vertebrae T5 vertebrae
iliac artery left cardiovascular system rib right 3 rib vertebrae T6 vertebrae
iliac artery right cardiovascular system rib right 4 rib vertebrae T7 vertebrae
iliac vena left cardiovascular system rib right 5 rib vertebrae T8 vertebrae
iliac vena right cardiovascular system rib right 6 rib vertebrae T9 vertebrae
iliopsoas left iliopsoas rib right 7 rib
We designed a CBIR benchmark relying on the TS dataset which is publicly available on Wasserthal et al. (2023). TS is a dataset comprising 1204 computed tomography (CT) volumes with 104 anatomical structure annotations. Since the anatomical regions presented in the original dataset include small structures we additionally mapped these small regions to classes with coarse labels, e.g., all the rib classes are mapped to a single class in the coarse label classes. The coarse label classes can provide insight into the retrieval of anatomical regions that are close to the target organ. LABEL:tab:_mapping_table shows the mapping of the TS original classes to the coarse classes. The query cases are sourced from the test split, while the train set serves as the database for searching. The search is assessed on the retrieval rate of 29 coarse anatomical structures and 104 original TS anatomical structures.

### v2

@@ Abstract
While content-based image retrieval (CBIR) has been extensively studied in natural image retrieval, its application to medical images presents ongoing challenges, primarily due to the 3D nature of medical images. Recent studies have shown the potential use of pre-trained vision embeddings for CBIR in the context of radiology image retrieval. However, a benchmark for the retrieval of 3D volumetric medical images is still lacking, hindering the ability to objectively evaluate and compare the efficiency of proposed CBIR approaches in medical imaging. In this study, we extend previous work and establish a benchmark for region-based and localized multi-organ retrieval using the TotalSegmentator dataset (TS) with detailed multi-organ annotations. We benchmark embeddings derived from pre-trained supervised models on medical images against embeddings derived from pre-trained unsupervised models on non-medical images for 29 coarse and 104 detailed anatomical structures in volume and region levels. For volumetric image retrieval, we adopt a late interaction re-ranking method inspired by text matching. We compare it against the original method proposed for volume and region retrieval and achieve a retrieval recall of 1.0 for diverse anatomical regions with a wide size range. The findings and methodologies presented in this paper provide insights and benchmarks for further development and evaluation of CBIR approaches in the context of medical imaging.
@@ 1 Introduction
In the realm of computer vision, content-based image retrieval (CBIR) has been the subject of extensive research for several decades (Dubey, 2021). CBIR systems typically utilize low-dimensional image representations stored in a database and subsequently retrieve similar images based on distance metrics or similarity measures of the image representations. Early approaches to CBIR involved manually crafting distinctive features, which led to a semantic gap, resulting in the loss of crucial image details due to the limitations of low-dimensional feature design (Dubey, 2021; Wang et al., 2022). However, recent studies in deep learning have redirected attention towards the creation of machine-generated discriminative feature spaces, effectively addressing and bridging this semantic gap (Qayyum et al., 2017). This shift has significantly enhanced the potential for more accurate and efficient CBIR methods (Dubey, 2021).
@@
While natural image retrieval has been extensively researched, the application of retrieval frameworks to medical images, particularly radiology images, presents ongoing challenges. CBIR offers numerous advantages for medical images. Radiologists can utilize CBIR to search for similar cases, enabling them to review the history, reports, patient diagnoses, and prognoses, thereby enhancing their decision-making process. In real-world use-cases, we often encounter huge anonymized and unannotated datasets available from different studies or institutions where the available meta-information, such as DICOM header data, has been removed or is inconsistent. Manually, searching for relevant images in such databases is extremely time-consuming. Moreover, the development of new tools and research in the medical field requires trustable dataset sources and therefore a reliable method for retrieving images, making CBIR an essential component in advancing computer-aided medical image analysis and diagnosis. One of the key challenges with applying standard CBIR techniques to medical images lies in the fact that algorithms developed for natural images are typically designed for 2D images, while medical images are often 3D volumes which adds a layer of complexity to the retrieval process.
@@
Recent studies have proposed and demonstrated the potential use of pre-trained vision embeddings for CBIR in the context of radiology image retrieval (Khun Jush et al., 2023; Abacha et al., 2023; Denner et al., 2024; Truong et al., 2023). However, these studies have primarily focused on 2D images (Denner et al., 2024) or specific pathologies or tasks (Abacha et al., 2023; Khun Jush et al., 2023; Truong et al., 2023), overlooking the presence of multiple organs in the volumetric images, which is a critical aspect of real-world scenarios. Large multi-organ medical image datasets can be leveraged to thoroughly evaluate the efficacy of the proposed methods, enabling a more comprehensive assessment of CBIR approaches for radiology images. Despite previous efforts, there is still no established benchmark available for comparing methods for the retrieval of 3D volumetric medical images. This absence of a benchmark impedes the ability to objectively evaluate and compare the efficiency of the proposed CBIR approaches in the context of medical imaging.
@@
Our previous work (Khun Jush et al., 2023) demonstrated the potential of utilizing pre-trained embeddings, originally trained on natural images, for various medical image retrieval tasks using the Medical Segmentation Decathlon Challenge (MSD) dataset (Antonelli et al., 2022). The approach is outlined in Figure 1. Building upon this, the current study extends the methodology proposed in Khun Jush et al. (2023) to establish a benchmark for anatomical region-based and localized multi-organ retrieval. While the focus of Khun Jush et al. (2023) was on evaluating the feasibility of using 2D embeddings and benchmarking different aggregation strategies of 2D information for 3D medical image retrieval within the context of the single-organ MSD dataset (Antonelli et al., 2022), it was observed that the single-organ labeling, hinders the evaluations for images containing multiple organs. The main objective of this study is to set a benchmark for organ retrieval at the localized level, which is particularly valuable in practical scenarios, such as when users zoom in on specific regions of interest to retrieve similar images of the precise organ under examination. To achieve this, we evaluate a count-based method in regions using the TotalSegmentator dataset (TS) (Wasserthal et al., 2023). TS dataset along with its detailed multi-organ annotations is a valuable resource for medical image analysis and research. This dataset provides comprehensive annotations for 104 organs or anatomical structures, which allow us to derive fine-grained retrieval tasks and comprehensively evaluate the proposed methods.
@@
  • We benchmarked pre-trained 2D embeddings trained supervised on medical images against self-supervised pre-trained embeddings trained on non-medical images for 3D radiology image retrieval. We utilize a count-based method to aggregate search results based on slice similarity to volume-level data retrieval.
  • We propose evaluation schemes based on the TotalSegmentator dataset Wang et al. (2022) for 29 aggregated coarse anatomical regions and all 104 original anatomical regions. Our proposed evaluation assesses the capabilities of a 3D image search system at different levels, including a fine-grained measure related to the localization of anatomical regions.
  • We adopted a late interaction re-ranking method originally used for text retrieval called ColBERT (Khattab and Zaharia, 2020) for volumetric image retrieval. For a 3D image query, this two-stage method generates a candidate 3D image result list utilizing a fast slice-wise similarity search and count-based aggregation. In the second stage, the full similarity information between all query and candidate slices is aggregated to determine re-ranking scores.
  • We benchmarked the proposed re-ranking method against the original method proposed in Khun Jush et al. (2023) for volume, region, and localized retrieval on 29 modified coarse anatomical regions and 104 original anatomical regions from TS dataset Wang et al. (2022).
@@ 2.1 Vector Database and Indexing
In the context of image search a database typically constitutes the central location where all the representations of the images, a.k.a. embeddings, and their metadata including annotations are stored. A query allows the user or the system to request specific images in various ways, e.g., by inputting a reference image or a textual description. The goal is to search the database for similar images that match the query. Similarly, in this study, the search process entails comparing a query image with images in the database to identify the most similar image using the cosine similarity of the embeddings. Throughout this process, we do not depend on any metadata information at any stage. Metadata-independence is an intended design choice and in contrast to widely used metadata-based image retrieval solutions that frequently lack the necessary specificity in real-world retrieval applications. In small sets, the similarity search is easy but with the growing size of the database, the complexity increases. Accuracy and speed are the key factors in search, thus, naive approaches typically fail in huge datasets.
@@
Indexing in the context of content-based image search involves creating a structured system that allows for efficient storage and retrieval of images based on their visual content. A flat index is the simplest form of indexing, where no modification is made to the vectors before they are used for search. In flat indexing, the query vector is compared to every other full-size vector in the database and their distances are calculated. The nearest k of the searched spaces is then returned as the k-nearest neighbors (kNN). While this method is the most accurate, it comes at the cost of significant search time (Aumüller et al., 2020). To improve search time, two approaches can be employed: reducing the vector size through dimensionality reduction, e.g., by reducing the number of bits representing each vector, or reducing the search scope by clustering or organizing vectors into tree structures based on similarity or distance. This results in the identification of an approximation of the true nearest neighbors, known as approximate nearest neighbor search (ANN) (Aumüller et al., 2020).
@@
There are several ANN methods available. In the context of content-based volumetric medical image retrieval, Khun Jush et al. (2023) compared Locality Sensitive Hashing (LSH) Charikar (2002) and Hierarchical Navigable Small World (HNSW) Malkov and Yashunin (2018) for indexing and search. LSH hashes data points in a way that similar data points are mapped to the same buckets with higher probabilities. This allows for a more efficient search for nearest neighbors by reducing the number of candidates to be examined. HNSW (Malkov and Yashunin, 2018) indexing organizes data into a hierarchical graph structure where each layer of the hierarchy has a lower resolution. The top layer connects data points directly, but the lower layers have fewer connections. The graph structure is designed to allow for efficient navigation during the search. Compared to LSH, HNSW typically enables faster search and requires less memory Taha et al. (2024). Based on findings in (Khun Jush et al., 2023) HSNW was chosen as the indexing method in the setting of this study due to speed advantages over LSH at a comparable recall. There are various index solutions available to store and search vectors. In this study, we used the Facebook AI Similarity Search (FAISS) package that enables fast similarity search (Johnson et al., 2019). The indexing process involves running the feature extractors on slices of each volumetric image and storing the output embeddings per slice. The produced representations are then added to the search index which is used later on for vector-similarity-based retrieval.
@@ 2.2 Feature Extractors
Self-supervised Models: We employed three self-supervised models pre-trained on ImageNet (Deng et al., 2009). DINOv1 (Caron et al., 2021), that demonstrated learning efficient image representations from unlabeled data using self-distillation. DINOv2 (Oquab et al., 2023), is built upon DINOv1 (Caron et al., 2021), and this model scales the pre-training process by combining an improved training dataset, patchwise objectives during training and introducing a new regularization technique, which gives rise to superior performance on segmentation tasks. DreamSim (Fu et al., 2023), built upon the foundation of DINOv1 (Caron et al., 2021), fine-tunes the model using synthetic data triplets specifically designed to be cognitively impenetrable with human judgments. For the self-supervised models, we used the best-performing backbone reported by the developers of the models.
Supervised Models: We included a SwinTransformer model (Liu et al., 2021) and a ResNet50 model (He et al., 2016) trained in a supervised manner using the RadImageNet dataset (Mei et al., 2022) that includes 5 million annotated 2D CT, MRI, and ultrasound images of musculoskeletal, neurologic, oncologic, gastrointestinal, endocrine, and pulmonary pathology. Furthermore, a ResNet50 model pre-trained on rendered images of fractal geometries was included based on (Kataoka et al., 2022). These training images are formula-derived, non-natural, and do not require any human annotation.
2.3 Dataset and Pre-processing
@@
adrenal gland left adrenal gland iliopsoas right iliopsoas rib right 11 rib
adrenal gland right adrenal gland inferior vena cava cardiovascular system rib right 12 rib
@@
gallbladder gallbladder rib left 2 rib vertebrae C5 vertebrae
gluteus maximus left gluteus muscles rib left 3 rib vertebrae C6 vertebrae
gluteus maximus right gluteus muscles rib left 4 rib vertebrae C7 vertebrae
gluteus medius left gluteus muscles rib left 5 rib vertebrae L1 vertebrae
gluteus medius right gluteus muscles rib left 6 rib vertebrae L2 vertebrae
gluteus minimus left gluteus muscles rib left 7 rib vertebrae L3 vertebrae
gluteus minimus right gluteus muscles rib left 8 rib vertebrae L4 vertebrae
heart atrium left cardiovascular system rib left 9 rib vertebrae L5 vertebrae
heart atrium right cardiovascular system rib left 10 rib vertebrae T1 vertebrae
heart myocardium cardiovascular system rib left 11 rib vertebrae T2 vertebrae
heart ventricle left cardiovascular system rib left 12 rib vertebrae T3 vertebrae
heart ventricle right cardiovascular system rib right 1 rib vertebrae T4 vertebrae
hip left hip rib right 2 rib vertebrae T5 vertebrae
hip right hip rib right 3 rib vertebrae T6 vertebrae
humerus left humerus rib right 4 rib vertebrae T7 vertebrae
humerus right humerus rib right 5 rib vertebrae T8 vertebrae
iliac artery left cardiovascular system rib right 6 rib vertebrae T9 vertebrae
iliac artery right cardiovascular system rib right 7 rib vertebrae T10 vertebrae
iliac vena left cardiovascular system rib right 8 rib vertebrae T11 vertebrae
iliac vena right cardiovascular system rib right 9 rib vertebrae T12 vertebrae
iliopsoas left iliopsoas rib right 10 rib
We designed a CBIR benchmark relying on the publicly available TotalSegmentator (TS) dataset Wasserthal et al. (2023), version 1. This dataset comprises in total of 1204 computed tomography (CT) volumes covering 104 anatomical structure annotations (TS, V1). The anatomical regions presented in the original dataset include several fine-grained sub-classes for which we considered an aggregation to a coarser common class as a reasonable measure, e.g., all the rib classes are mapped to a single class ‘rib’. The coarse organ labels can help identify similarities and potential mismatches between neighboring anatomical regions, providing valuable insights into the proximity information of the target organ. LABEL:tab:_mapping_table shows the mapping of the original TS classes to the coarse aggregated classes. For the sake of reproducibility, the query cases are sourced from the original TS test split, while the cases contained in the original TS train and validation set serve as the database for searching. The search is assessed on the retrieval rate of 29 coarse anatomical structures and 104 original TS anatomical structures.